. Pat and Kim bought the same item. Pat said he paid 20% less than the list price. Kim said that if she paid 25% more for the item, she would have paid the list price. Who paid the least? 


1
Expert's answer
2021-03-09T05:33:20-0500

Solution:

Let the list price be $ 100.

Case I: Pat bought it for 20% less, thus he paid =(100-20)= $ 80 (As discount is calculated on the list price)

Case II: Let Kim bought it for $x.

The equation formed: x+0.25x =100

1.25x=100

x=100/1.25

x=80

Thus Kim bought it for $80, which is the same price Pat bought it for. So, they bought it for the same price.